OpenStudy (anonymous):

Use the formula for slope to find the missing coordinates. 1.) The points (2,4) and (5,_) lie on a line whose slope is 7/3.

OpenStudy (anonymous):

(y-4)/(5-2) = 7/3

OpenStudy (anonymous):

the formula is \[y2-y1/x2-x1\]

OpenStudy (anonymous):

yes if you knew it you could find answer.

OpenStudy (anonymous):

i dont know how to find the answer

OpenStudy (anonymous):

solve the expression i gave you. you know how to solve it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then solve it and find the value y it's your missing term.

OpenStudy (anonymous):

m = (y2-y1)/(x2-x1) (x-4)/(5-2)= 7/3 (x-4) = 3*7/3 x-4=7 x=11 .. (5,11)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

as question (1/4,-n) and (3/4,n)

OpenStudy (anonymous):

how to solve

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

as question (1/4,-n) and (3/4,n)

OpenStudy (anonymous):

how to solve this

OpenStudy (anonymous):

you want to calculate slope or what?

OpenStudy (anonymous):

calculate

OpenStudy (anonymous):

In a equation

OpenStudy (anonymous):

you gave me two points (1/4,-n) and (3/4,n) but don't specify what to do what i have to solve?

OpenStudy (anonymous):

plzz solve

OpenStudy (anonymous):

ha ha i don't get your question please ask as a new, someone get it and solve it.

OpenStudy (anonymous):

(1/4,-n) and (3/4,n) how to find thier slope

OpenStudy (anonymous):

same as m = (n+n)/(3/4 -1/4)=4n